A company manufactures bulbs. The probability of getting a defective bulb is 0.045. A sample of 100 bulbs were selected. Use Poisson approximation to binomial distribution to find the probability of finding at most 3 non-defective bulbs


1
Expert's answer
2021-07-16T01:47:23-0400

The Poisson distribution is

"P\left(X=x\right)=\frac{e^{-\lambda }\lambda ^x}{x!}"

Let "x" denote the non - defective fuse

"P=1-0.045=0.955"

"n=100"

"mean=\lambda =np=100\times0.955=95.5"

"P\left(atmost\:3\:non-defective\:fuses\right)= \sum_{x=0}^3 \frac{e^{-95.5}95.5^x}{x!}"

"P\left(atmost\:3\:non-defective\:fuses\right)=" "0.01528"